OpenStudy (anonymous):

find the general solutions of : tan x+ cot 2x = 0

OpenStudy (anonymous):

fisrt change in terms of sin and cos and then simplify fuirther

OpenStudy (anonymous):

tan x+ cot 2x = 0 sinx/cosx + cos2x/sin2x =0 sinxsin2x+cosxcos2x= 0 cos(2x-x)=0 or cosx =0....and further..

OpenStudy (mathmale):

Note that there is a trig identity for cot 2x. You might want to look up "trig identities" and look for it. Alternatively, you could re-write cot 2x as cos2x/sin2x, as "matricked" has done.

OpenStudy (anonymous):

even i got cos x = 0 but the general solution given is n (pi)- pi/2 .. which i am not getting @matricked @mathmale

OpenStudy (anonymous):

if cos x =0 then x= odd multiples of (pi/2)

OpenStudy (anonymous):

x= (2n-1)(pi/2) = n(pi) - (pi/2)

OpenStudy (mathmale):

A solution involving n, such as this one, stems from the periodicity of the sine and cosine functions. The cosine function repeats itself every 2Pi radians. cos x = 0 at x = Pi/2; this is true again at 2Pi + Pi/2, at 4Pi + Pi/2, and so on; that's where the n2Pi comes from. Note that cos x is 0 also at x =3Pi/2; the same pattern happens here also.